# Automatic Combination of Sample Selection Strategies for Few-Shot Learning

### 05 Feb 2024 v1

@@ 4 Experiments
Datasets. The experiments are conducted on image and text classification datasets composed of different tasks with different number of classes. For the image datasets, we use 8 randomly chosen datasets from the Meta-Album (Ullah et al., 2022), specifically LR_AM.DOG and LR_AM.AWA for animal classification, VCL.APL for plane classification, HUM_ACT.ACT_410 for classification of human actions, MNF.TEX_DTD for classification of texture, MCR.PRT and MCR.PNU for classification of microscope images, and PLT.FLW for flower classification. In addition, we use 6 widely-adopted text datasets: 20 News Group (Lang, 1995) and News Category (Misra, 2022) for news category classification and ATIS (Hemphill et al., 1990), Facebook (Schuster et al., 2019), HWU-64 (Liu et al., 2021) and SNIPS (Coucke et al., 2018) for intent classification.
@@
Few-shot Learning Approaches. The evaluation is done in a 5-way 5-shot setting using the gradient few-shot learning on both the image and text datasets and the in-context learning on the text datasets. For gradient few-shot learning, we use meta-learning approaches, specifically Prototypical Networks (Snell et al., 2017) and MAML (Finn et al., 2017), and Few-Shot Fine-Tuning (Chen et al., 2019). For in-context learning, we use the Mistral-7B (Jiang et al., 2023) and Zephyr-7B (Tunstall et al., 2023) models.

### 17 Apr 2026 v2

@@ 4 Experiments
Datasets. We focus on representative few-shot learning datasets composed of different tasks with different numbers of classes and with different characteristics, such as class imbalance or amount of noisy samples or labels. Specifically, we use 6 widely-adopted text datasets, composed of tasks with different number of classes and characteristics, specifically: 20 News Group (Lang, 1995) and News Category (Misra, 2022) for news category classification and ATIS (Hemphill et al., 1990), Facebook (Schuster et al., 2019), HWU-64 (Liu et al., 2021) and SNIPS (Coucke et al., 2018) for intent classification. For each dataset, we select a subset of up to 200 labelled samples per class to simulate the limited data setting if needed (the impact of smaller dataset sizes is analysed in Appendix I). Furthermore, to evaluate the generalisability, we repeat the experiments on image data as well (see Appendix E for details).
@@
Few-shot Learning Approaches. The evaluation is done in a 5-way 5-shot setting using in-context learning on the text datasets and the gradient few-shot learning on both the text and image datasets. For in-context learning, we use the Mistral-7B (Jiang et al., 2023), Zephyr-7B (Tunstall et al., 2023), LLaMA-3.1-8B (Grattafiori et al., 2024), Gemma-3-4B (Team, 2025) and Qwen-2.5-7B (Team, 2024) instruction-tuned large language models. For gradient few-shot learning, we use representative and widely-used few-shot learning approaches, specifically 1) meta-learning approaches Prototypical Networks (Snell et al., 2017) and MAML (Finn et al., 2017), and 2) Few-Shot Fine-Tuning (Chen et al., 2019).